# Interaction-Induced Dimensional Crossover through Full 3D to 1D

### v2

@@ Abstract
The exploration of dimensional crossover carries profound fundamental significance, serving as a crucial bridge in comprehending the remarkable disparities observed in transitional phenomena across the two distinct dimensions of a physical system. The prevalent strategy for manipulating the dimensionality involves meticulously controlling the external trapping geometry, thereby restricting the degrees of freedom of the kinetic energy from three-dimensional (3D) to lower-dimensional spaces, while maintaining the 3D nature of the interaction energy degrees of freedom. The aim of this work is to introduce an innovative scenario to achieve dimensional crossover, characterized by lower-D nature of both the kinetic and the interaction energy degrees of freedom. To accomplish this objective, we delve deeply into the realm of a 2D optically trapped Bose gas, focusing specifically on its finite-range interaction. Our emphasis lies in exploring the lattice-induced dimensional crossover from full 3D to 1D in both kinetic and interaction terms. Utilizing the functional path integral method, we derive the equation of states of the model system, encompassing crucial quantities such as the ground state energy and quantum depletion. These equations enable us to analyze the combined effects of finite range interaction and an optical lattice on quantum fluctuations of the BEC system. Notably, our analytical findings reconcile the Lee-Huang-Yang (LHY) correction to the ground state energy in 3D and Lieb-Liniger (LL) ones in 1D limit, thereby providing fresh insights into the intriguing disparities between LHY and LL corrections.
@@ I INTRODUCTION
Currently, there exists significant and ongoing interests in exploring how the dimensionality of a physical system impacts its fundamental nature and properties [1]. The motivation behind this exploration is twofold. Firstly, lower dimensions often amplify quantum and thermal fluctuations within the model system, leading to a rich array of genuine quantum effects as dimensionality decreases [2, 3, 4]. A typical example consists in the interacting bosons, characterized by the superfluids in three dimension (3D) [5], Berezinski-Kosterlitz-Thouless (BKT) topological phase transitions in 2D [6, 7] and Tomonaga-Luttinger liquids in 1D [8] repectively. Secondly, cutting-edge experimental techniques provide remarkable precision in controlling key parameters of the physical system, such as interaction strength and confining potentials [1]. In particular, the systm’s dimensionality can be set freely rather than being fixed once for all. In the context of ultracold gases, highlights in 1D and 2D include the observation of bosonic fermionization into the Tonks-Girardeau (TG) state [9, 10, 11], the expansion dynamics [12] and the Super-Tonks-Girardeau quench [13] of dipolar bosons. The stark contrast between the aforementioned properties in lower dimensions and their 3D counterparts underscores the importance of exploring the dimensional crossover between these two distinct dimensions.
@@
Along this research line, considerable efforts have been dedicated to gaining a profound understanding of the dimensional crossovers, ranging from a higher to a lower dimension, within the realm of ultracold quantum systems. For instance, Refs. [14, 15, 16, 17, 18] have delved into the dimensional crossover exhibited by Bose gas, as well as the counterpart in Fermi gas [19, 20, 21, 22]. Through the meticulous manipulation of the transverse $z$-direction, Refs. [23, 24, 25, 26] have successfully achieved a dimensional crossover from 3D to 2D. Furthermore, Refs. [27, 28] have demonstrated that the introduction of an external trapping potential gives rise to a crossover from 3D to 1D. It is noteworthy that recent theoretical and experimental [29, 30, 31] investigations have revealed the dimensional crossover of strongly-interacting bosons from 2D to 1D, highlighting the rich and complex nature of these systems across different dimensions.
@@
Dimensional crossovers are distinguished by their hierarchical access to novel energy and length scales. Leveraging this understanding, the predominant approach in realizing dimensional crossover during the last decade is that a physical system becomes quasi-low dimensional when energetic restriction to freeze excitations in one or two dimensions is reached, which is referred as to the kinetic-induced low-dimensional physical system [27, 28, 29, 31, 30, 32]. It is important to note that the interactions in such low-dimensional systems are often treated analogously to their 3D counterparts. However, in the following, we introduce an alternative scenario based on a 2D optically-trapped Bose gas with weak interactions and specifically engineered finite-range interatomic interactions. As we will demonstrate, considering finite-range interactions offers a novel perspective on low-dimensional systems, characterized by both their interactions and kinetic energy being inherently low-dimensional.
@@
The second impetus for this work stems from recent experimental advancements in realizing ultracold quantum gases with finite-range interatomic interaction [33, 34]. Compare with the quantum gases exhibiting s-wave interactions, ultracold quantum gases with finite-range interatomic interactions possess crucial novel features, as outlined below: the equation of state (EOS) for s-wave interacting quantum gas exhibits a profound universality, stemming from the fact that it can be precisely characterized by a single parameter: the s-wave scattering length $a_{\text{s}}$. This parameter serves as a versatile tool, encapsulating not just the fundamental nature of two-body interactions but also the intricate many-body physics [35, 36] of the system. However, the introduction of finite-range atomic interactions disrupts this universality, giving rise to nonuniversal effects within the EOS [37, 38, 39, 40, 41]. At the mean-field level, the Gross-Pitaevskii equation undergoes modifications [42, 43, 38, 44] to incorporate these nonuniversal effects. The modified Gross-Pitaevskii equation provides a comprehensive framework for elucidating the physical behavior of nonuniform condensates, effectively capturing deviations from the idealized s-wave interaction scenario. Beyond the mean-field approximation, recent studies have broadened the thermodynamic analysis of pure two-dimensional and three-dimensional uniform Bose gases [45, 46] to incorporate finite-range interactions, reaching the sophisticated Gaussian level of approximation. We mention that our previous work [47] has investigated finite-range-interaction induced EOS along dimensional crossover from 3D to 2D. These works provide a deeper comprehension of the rich and intricate physics that arises in the presence of non-ideal interactions.
@@
In this work, we have delved into the intricate phenomenon of dimensional crossover exhibited by Bose gases with finite-range interactions confined within optical lattices. Employing the finite-temperature functional path integral framework and the tight-binding approximation, we have embarked on a comprehensive exploration of the ground-state properties of these Bose gases in a two-dimensional optical lattice setting. Our findings have culminated in the derivation of a beyond-mean-field equation of state for the Bose system, one that encompasses the influence of various parameters including optical lattice configurations, effective ranges, and scattering lengths. This equation of state yields generalized corrections to the Lee-Huang-Yang (LHY) energy in 3D and the Lieb-Liniger (LL) solution in 1D. Furthermore, by manipulating the effective range, we have visually analyzed the subsequent influence of finite-range interactions on Gaussian quantum fluctuations, offering a deeper insight into their pivotal role in dimensional crossover phenomena.
@@
The structure of this paper is outlined as follows. In Sec. II, we introduce the action functional of the model system and demonstrate the fundamental terminology of the finite-temperature functional path integral. In Sec. III, we calculate the ground state energy and quantum depletion of the system through thermodynamic relationships and regularization method. Additionally, we analyze the asymptotic behavior of these two physical quantities under both the 1D and 3D limits. Sec. IV summarizes our work, and we also consider the potential experimental conditions for realizing our proposed scenario.
@@ II Interaction-Induced Dimensional Crossover
A common strategy to achieving the low-D system in an ultra-cold quantum gas is to add an optical lattice to the microscopic Hamiltonian [27, 23]. To illustrate this, let’s consider the dimensional crossover from 3D to quasi-1D by introducing a 2D optical lattice as an example. The dispersion of single-particle energy in the model system is given by $\epsilon_{\mathbf{k}}^{0}=\hbar^{2}k_{z}^{2}/2m+J\left[2-\cos k_{x}d-\cos k_{y}d\right]$ with $J$ being the tunneling rate between two optical wells. When the energetic constraints freeze excitations along $x$- and $y$- directions, the system effectively becomes quasi-1D with the energy dispersion simplified into $\epsilon_{\mathbf{k}}^{0}=\hbar^{2}k_{z}^{2}/2m$ as $J\rightarrow 0$. Transitioning into the quasi-1D regime, the coupling constant is renormalized as $\tilde{g}=C^{2}g_{0}$ with $g_{0}$ being the 3D coupling constant and $C$ being the renormalized constant [27]. We refer it as to the kinetic-induced 1D physical system, distinguished by its 1D kinetic energy and 3D interaction energy.
@@
with $g_{0}=4\pi\hbar^{2}a_{s}/m$ and $g_{2}=2\pi\hbar^{2}a_{s}^{2}r_{s}/m$ being the s-wave scattering and finite-range coupling constant. Here, $a_{s}$ is the s-wave scattering length and $r_{s}$ represts the effective range of interatomic interaction potential [49]. With the introduction of a 2D optical lattice, the interaction energy corresponding to Eq. (1) is proportional to be $\tilde{g}_{2}\left[k_{z}^{2}+J_{1}\left(2-\cos k_{x}d-\cos k_{y}d\right)\right]$. By changing $J_{1}\rightarrow 0$, one can reach the regime of 1D interaction with the form of $\tilde{g}_{2}k_{z}^{2}$. In what follows, we focus on the equation of state of the interacting Bose gas trapped into a 2D optical lattice and demonstrate the interaction-induced dimensional crossover through full 3D to qusi-1D.
@@
We confine our discussion to the scenario where the lattice depth $t$ is substantial enough to render the interwell barriers significantly higher than the chemical potential $\mu$ [27, 23, 28, 56]. Under such conditions, numerous isolated wells are formed, giving rise to an orderly array of condensates. Simultaneously, due to quantum tunneling, the overlap between the wave functions of two continuous wells are still sufficient to ensure full coherence. Based on this assumption, we refer to the density of the condensate as $n_{0}$ and ignore the phase transition of the Mott insulator.
@@
First, the functional of Eq. (5) has two importances with respect to the functional characterizing a trapped Bose gas in the absence of optical confinement: (i) the kinetic energy in the first line of the functional of Eq. (5) exhibits a periodic dependence as $\varepsilon^{0}_{\mathbf{k}}=J(2-\cos k_{x}d-\cos k_{y}d)+\hbar^{2}k_{z}^{2}/2m$. In the 3D limit when the lattice strength is weak, the kinetic energy can be rewritten as $\varepsilon^{0}_{\mathbf{k}}=\hbar^{2}(k_{x}^{2}+k_{y}^{2})/2m^{*}+\hbar^{2}k_{z}^{2}/2m$ with $m^{*}=\hbar^{2}/Jd^{2}$. With the increase of lattice strength, the value of $J$ will decay exponentially. As a result, the kinetic energy experiences a dimensional crossover from 3D to 1D and becomes to be an effective 1D one of $\varepsilon^{0}_{\mathbf{k}}=\hbar^{2}k_{z}^{2}/2m$. (ii) The s-wave scattering coupling constant in the second line of the functional of Eq. (5) is renormalized as $\tilde{g}_{0}=C^{2}g_{0}$ with $C=\sqrt{\pi/2}t^{1/4}e^{-1/4\sqrt{t}}$.
Second, the finite-range interaction term in the third line of the functional of Eq. (5) exhibits the dynamical 1D features of particle motion, and a dimensional crossover from 3D to 1D emerges in the behavior of the interaction energy when the energetic restriction to freeze anxial excitations is reached.
(i) For $8J/\mu\gg 1$, the system exhibits a distinct anisotropic 3D behavior, and the finite-range effective coupling constant takes form of $\frac{\tilde{g}_{2}\tilde{q}^{2}}{2}=\frac{g_{2}C^{2}m}{\hbar^{2}}\times\left[\frac{\hbar^{2}(k_{x}^{2}+k^{2}_{y})}{2m(2/J_{1}d^{2})}+\frac{\hbar^{2}k_{z}^{2}}{2m}\right]$.
@@
Now, we are ready to calculate the analytical expressions of the equation of state of model system along the dimensional crossover. In this end, by applying the mean field plus Gaussian (one loop) approximation [45] to Eq. (5) and writing $\sum_{\mathbf{k},n}\psi_{\mathbf{k},n}=\sqrt{n_{0}}+\sum_{\mathbf{k},n\neq 0}\phi_{\mathbf{k},n}$, we find that the Gaussian contribution of quantum fluctuation is described by
@@
where the chemical potential being $\mu=\tilde{g}_{0}n_{0}$ and $\tilde{V}\left(\mathbf{k}\right)=\tilde{g}_{0}+\tilde{g}_{2}\left[k_{z}^{2}+J_{1}\left(2-\cos k_{x}d-\cos k_{y}d\right)\right]$. By integrating on the Boson field $\phi_{\mathbf{k},n}$ and $\phi_{\mathbf{k},n}^{*}$, after that the sum over bosonic Matsubara frequencies gives [57] the Gaussian grand potential
$\Omega_{g}=\frac{E_{\mathbf{k}}}{2}+\frac{1}{\beta}\ln\left(1-e^{-\beta E_{\mathbf{k}}}\right),$ (8)

### v3

@@ Abstract
The exploration of dimensional crossover carries profound fundamental significance, serving as a crucial bridge in comprehending the remarkable disparities observed in transitional phenomena across the two distinct dimensions of a physical system. The prevalent strategy for manipulating the dimensionality involves meticulously controlling the external trapping geometry, thereby restricting the degrees of freedom of the kinetic energy from three-dimensional (3D) to lower-dimensional spaces, while maintaining the 3D nature of the interaction energy degrees of freedom. The aim of this work is to introduce an innovative scenario to achieve dimensional crossover, characterized by lower-D nature of both the kinetic and the interaction energy degrees of freedom. To accomplish this objective, we delve deeply into the realm of a 2D optically trapped Bose gas, focusing specifically on its finite-range interaction. Our emphasis lies in exploring the lattice-induced dimensional crossover from full 3D to 1D in both kinetic and interaction terms. Utilizing the functional path integral method, we derive the equation of states of the model system, encompassing crucial quantities, such as the ground state energy and quantum depletion. These equations enable us to analyze the combined effects of finite range interaction and an optical lattice on quantum fluctuations of the BEC system. Notably, our analytical findings reconcile the Lee-Huang-Yang (LHY) correction to the ground state energy in 3D and Lieb-Liniger (LL) ones in 1D limit, thereby providing fresh insights into the intriguing disparities between LHY and LL corrections.
@@ I INTRODUCTION
Currently, there are significant and ongoing interests in exploring how the dimensionality of a physical system impacts its fundamental nature and properties [1]. The motivation behind this exploration is twofold. Firstly, lower dimensions often amplify quantum and thermal fluctuations within the model system, leading to a rich array of genuine quantum effects as dimensionality decreases [2, 3, 4]. A typical example consists of the interacting bosons, characterized by the superfluids in three dimension (3D) [5], Berezinski-Kosterlitz-Thouless (BKT) topological phase transitions in 2D [6, 7] and Tomonaga-Luttinger liquids in 1D [8] repectively. Secondly, cutting-edge experimental techniques provide remarkable precision in controlling key parameters of the physical system, such as interaction strength and confining potentials [1]. In particular, the system’s dimensionality can be set freely rather than being fixed once for all. In the context of ultracold gases, highlights in 1D and 2D include the observation of bosonic fermionization into the Tonks-Girardeau (TG) state [9, 10, 11], the expansion dynamics [12] and the Super-Tonks-Girardeau quench [13] of dipolar bosons. The stark contrast between the aforementioned properties in lower dimensions and their 3D counterparts underscores the importance of exploring the dimensional crossover between these two distinct dimensions.
@@
Along this research line, considerable efforts have been dedicated to gaining a profound understanding of the dimensional crossovers, ranging from a higher to a lower dimension, within the realm of ultracold quantum systems. For instance, Refs. [14, 15, 16, 17, 18] have delved into the dimensional crossover exhibited by Bose gas, as well as the counterpart in Fermi gas [19, 20, 21, 22]. Through the meticulous manipulation of the transverse $z$-direction, Refs. [23, 24, 25, 26, 27] have successfully achieved a dimensional crossover from 3D to 2D. Furthermore, Refs. [28, 29] have demonstrated that the introduction of an external trapping potential gives rise to a crossover from 3D to 1D. It is noteworthy that recent theoretical and experimental [30, 31, 32] investigations have revealed the dimensional crossover of strongly-interacting bosons from 2D to 1D, highlighting the rich and complex nature of these systems across different dimensions.
@@
Dimensional crossovers are distinguished by their hierarchical access to novel energy and length scales. Leveraging this understanding, the predominant approach to realizing dimensional crossover during the last decade is that a physical system becomes quasi-low dimensional when energetic restriction to freeze excitations in one or two dimensions is reached, which is referred to as the kinetic-induced low-dimensional physical system [28, 29, 30, 32, 31, 33]. It is important to note that the interactions in such low-dimensional systems are often treated analogously to their 3D counterparts. However, in the following, we introduce an alternative scenario based on a 2D optically-trapped Bose gas with weak interactions and specifically engineered finite-range interatomic interactions. As we will demonstrate, considering finite-range interactions offers a novel perspective on low-dimensional systems, characterized by both their interactions and kinetic energy being inherently low-dimensional.
@@
The second impetus for this work stems from recent experimental advancements in realizing ultracold quantum gases with finite-range interatomic interaction [34, 35]. Compared with the quantum gases exhibiting s-wave interactions, ultracold quantum gases with finite-range interatomic interactions possess crucial novel features, as outlined below: The equation of state (EOS) for s-wave interacting quantum gas exhibits a profound universality, stemming from the fact that it can be precisely characterized by a single parameter: the s-wave scattering length $a_{\text{s}}$. This parameter serves as a versatile tool, encapsulating not just the fundamental nature of two-body interactions but also the intricate many-body physics [36, 37] of the system. However, the introduction of finite-range atomic interactions disrupts this universality, giving rise to nonuniversal effects within the EOS [38, 39, 40, 41, 42]. At the mean-field level, the Gross-Pitaevskii equation undergoes modifications [43, 44, 39, 45] to incorporate these nonuniversal effects. The modified Gross-Pitaevskii equation provides a comprehensive framework for elucidating the physical behavior of nonuniform condensates, effectively capturing deviations from the idealized s-wave interaction scenario. Beyond the mean-field approximation, recent studies have broadened the thermodynamic analysis of pure two-dimensional and three-dimensional uniform Bose gases [46, 47] to incorporate finite-range interactions, reaching the sophisticated Gaussian level of approximation. We mention that our previous work [48] has investigated finite-range-interaction induced EOS along a dimensional crossover from 3D to 2D. These works provide a deeper comprehension of the rich and intricate physics that arises in the presence of non-ideal interactions.
@@
In this work, we have delved into the intricate phenomenon of dimensional crossover exhibited by Bose gases with finite-range interactions confined within optical lattices. Employing the finite-temperature functional path integral framework and the tight-binding approximation, we have embarked on a comprehensive exploration of the ground-state properties of these Bose gases in a two-dimensional optical lattice setting. Our findings have culminated in the derivation of a beyond-mean-field equation of state for the Bose system, one that encompasses the influence of various parameters, including optical lattice configurations, effective ranges, and scattering lengths. This equation of state yields generalized corrections to the Lee-Huang-Yang (LHY) energy in 3D and the Lieb-Liniger (LL) solution in 1D. Furthermore, by manipulating the effective range, we have visually analyzed the subsequent influence of finite-range interactions on Gaussian quantum fluctuations, offering a deeper insight into their pivotal role in dimensional crossover phenomena.
@@
The structure of this paper is outlined as follows: In Sec. II, we introduce the action functional of the model system and demonstrate the fundamental terminology of the finite-temperature functional path integral. In Sec. III, we calculate the ground state energy and quantum depletion of the system through thermodynamic relationships and the regularization method. Additionally, we analyze the asymptotic behavior of these two physical quantities under both the 1D and 3D limits. Sec. IV summarizes our work, and we also consider the potential experimental conditions for realizing our proposed scenario.
@@ II Interaction-Induced Dimensional Crossover
A common strategy for achieving the low-D system in an ultra-cold quantum gas is to add an optical lattice to the microscopic Hamiltonian [28, 23]. To illustrate this, let’s consider the dimensional crossover from 3D to quasi-1D by introducing a 2D optical lattice as an example. The dispersion of single-particle energy in the model system is given by $\epsilon_{\mathbf{k}}^{0}=\hbar^{2}k_{z}^{2}/2m+J\left[2-\cos k_{x}d-\cos k_{y}d\right]$ with $J$ being the tunneling rate between two optical wells. When the energetic constraints freeze excitations along $x$- and $y$- directions, the system effectively becomes quasi-1D with the energy dispersion simplified into $\epsilon_{\mathbf{k}}^{0}=\hbar^{2}k_{z}^{2}/2m$ as $J\rightarrow 0$. Transitioning into the quasi-1D regime, the coupling constant is renormalized as $\tilde{g}=C^{2}g_{0}$ with $g_{0}$ being the 3D coupling constant and $C$ being the renormalized constant [28]. We refer to it as the kinetic-induced 1D physical system, distinguished by its 1D kinetic energy and 3D interaction energy.
@@
with $g_{0}=4\pi\hbar^{2}a_{s}/m$ and $g_{2}=2\pi\hbar^{2}a_{s}^{2}r_{s}/m$ being the s-wave scattering and finite-range coupling constants. Here, $a_{s}$ is the s-wave scattering length and $r_{s}$ represents the effective range of interatomic interaction potential [50]. With the introduction of a 2D optical lattice, the interaction energy corresponding to Eq. (1) is proportional to be $\tilde{g}_{2}\left[k_{z}^{2}+J_{1}\left(2-\cos k_{x}d-\cos k_{y}d\right)\right]$. By changing $J_{1}\rightarrow 0$, one can reach the regime of 1D interaction in the form of $\tilde{g}_{2}k_{z}^{2}$. In what follows, we focus on the equation of state of the interacting Bose gas trapped in a 2D optical lattice and demonstrate the interaction-induced dimensional crossover through full 3D to qusi-1D.
@@
We proceed to provide approximate estimates of the parameter ranges where the tight-binding approximation remains valid. For this purpose, we adopt the commonly encountered experimental parameters for an optically trapped Bose gas reported in Ref. [57]. It is crucial to note that the tight-binding approximation holds under specific conditions. Primarily, the lattice depth $t$ in Eq. (4) must be relatively large $\left(t\geq 5\right)$ to ensure that the interband gap $E_{\text{gap}}$ exceeds the chemical potential $\mu$ [28, 23, 29, 58]. Secondly, under such conditions, numerous isolated wells are formed, giving rise to an orderly array of condensates. Simultaneously, due to quantum tunneling, the overlap between the wave functions of two continuous wells is still sufficient to ensure full coherence. Additionally, the typical detailed parameters read as follows: The recoil energy is $E_{R}\approx h\times 3.33$ kHz with $h$ being the Plank constant and the chemical potential of gas is $\mu\approx\tilde{g}_{0}n_{0}\approx h\times 400$ kHz. In the case of $t\approx 5$, we can estimate the parameters of $J/E_{R}\approx 0.18$ based on Ref. [59]. Consequently, the dimensionless parameters used in the figures of this work are as follows: $s=2J/\tilde{g}_{0}n_{0}\approx 3.6$, indicating that our model system exhibits 3D-like. Furthermore, as shown in  [57], the optically-trapped Bose gas exhibits complete superfluidity below the critical lattice height $t_{c}\approx 13$ corresponding to $J$ and $J_{1}$ being nearly zero [59]. We, therefore, conclude that the tight-binding approximation can be regarded as valid within the range of $135<t<$ 13, corresponding to $40<s<$ 4 as depicted in Fig. 1. Based on this assumption, we refer to the density of the condensate as $n_{0}$ and ignore the phase transition of the Mott insulator.
@@
Firstly, the functional of Eq. (5) has two importances with respect to the functional characterizing a trapped Bose gas in the absence of optical confinement: The kinetic energy in the first line of the functional of Eq. (5) exhibits a periodic dependence as $\varepsilon^{0}_{\mathbf{k}}=J(2-\cos k_{x}d-\cos k_{y}d)+\hbar^{2}k_{z}^{2}/2m$. In the 3D limit when the lattice strength is weak, the kinetic energy can be rewritten as $\varepsilon^{0}_{\mathbf{k}}=\hbar^{2}(k_{x}^{2}+k_{y}^{2})/2m^{*}+\hbar^{2}k_{z}^{2}/2m$ with $m^{*}=\hbar^{2}/Jd^{2}$. With the increase of lattice strength, the value of $J$ will decay exponentially. As a result, the kinetic energy experiences a dimensional crossover from 3D to 1D and becomes to be an effective 1D one of $\varepsilon^{0}_{\mathbf{k}}=\hbar^{2}k_{z}^{2}/2m$. (ii) The s-wave scattering coupling constant in the second line of the functional of Eq. (5) is renormalized as $\tilde{g}_{0}=C^{2}g_{0}$ with $C=\sqrt{\pi/2}t^{1/4}e^{-1/4\sqrt{t}}$.
Secondly, the finite-range interaction term in the third line of the functional of Eq. (5) exhibits the dynamical 1D features of particle motion, and a dimensional crossover from 3D to 1D emerges in the behavior of the interaction energy when the energetic restriction to freeze anxial excitations is reached.
(i) For $8J/\mu\gg 1$, the system exhibits a distinct anisotropic 3D behavior, and the finite-range effective coupling constant takes the form of $\frac{\tilde{g}_{2}\tilde{q}^{2}}{2}=\frac{g_{2}C^{2}m}{\hbar^{2}}\times\left[\frac{\hbar^{2}(k_{x}^{2}+k^{2}_{y})}{2m(2/J_{1}d^{2})}+\frac{\hbar^{2}k_{z}^{2}}{2m}\right]$.
@@
Now, we are ready to calculate the analytical expressions of the equation of state of the model system along the dimensional crossover. In this end, by applying the mean field plus Gaussian (one loop) approximation [46] to Eq. (5) and writing $\sum_{\mathbf{k},n}\psi_{\mathbf{k},n}=\sqrt{n_{0}}+\sum_{\mathbf{k},n\neq 0}\phi_{\mathbf{k},n}$, we find that the Gaussian contribution of quantum fluctuation is described by
@@
where the chemical potential being $\mu=\tilde{g}_{0}n_{0}$ and $\tilde{V}\left(\mathbf{k}\right)=\tilde{g}_{0}+\tilde{g}_{2}\left[k_{z}^{2}+J_{1}\left(2-\cos k_{x}d-\cos k_{y}d\right)\right]$. By integrating on the Boson fields $\phi_{\mathbf{k},n}$ and $\phi_{\mathbf{k},n}^{*}$, the sum over bosonic Matsubara frequencies gives [60] the Gaussian grand potential
$\Omega_{g}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{\mathbf{k}}}\frac{E_{\mathbf{k}}}{2}+\frac{1}{\beta}\ln\left(1-e^{-\beta E_{\mathbf{k}}}\right),$ (8)